OpenStudy (anonymous):

In a thermometer calibration test 0.31 OHM are registered when the temperature is 422 (K), according to R=R0*e^(A*((1/T)-(1/T0))) where A is a constant, R0 is in OHMs, and T0 in (K). If a R0=2.2 OHM for T0=310 (K), calculate R for 390 (°F).

OpenStudy (anonymous):

First, you need to find the value of A. Use the initial conditions along with the measured readings to find A. R=Ro*e^A*(1/T-1/To) so A = [ln(R/Ro)]/(1/T-1/To) = [ln(.31/2.2)]/(1/422-1/310) to find R for 390 F, convert F to Kelvin... 390 F = 472 K then used this value of T along with the value of A you calculated. so R=Ro*e^A*(1/T-1/To) =2.2*e^2289*(1/472-1/310)

OpenStudy (anonymous):

thank you very much, I thought the first condition was R=R0*e^A*(1/T-1/422), the second R=0.22*e^A*(1/T-1/310) and the last R=Ro*e^A*(1/T-1/472), because only T0 is the registered temperature, I'll ask about both ideas, maybe the problem was not worded nicely